Monday 04/07/14

9:30a – 11:30p


  • lab meeting (see protected notes)
  • my journal club presentation (see google-docs presentation)
  • initial planning meeting with Steven — need joint write-up, goal to write by Wed


  • finalizing write-up 1 for XZ with Jeff.
  • working on rewriting write-up 2 with team.

Chromatin imaging

  • samples incubated an extra 24 hours.
  • no staining! (in either 647 or 750)
  • hypotheses:
    1. maybe the humidity chamber condensation which hydrates the rubber cement dilutes away the probe?
    • try dry hybe comparison tonight
      1. maybe something is really wrong with new hybe dilution mix
    • I think this mix worked sorta the first time the 750 failed last week.
  • took bead images (currently on STORM2 hard-drive, need to copy over).
  • imaging sample E02

New Staining

  • F07-P1 F06-P3 4:4. .33 A647 and .66 A750
  • new E04-P1 2.5 uL to .33uL A647
  • new E05-P1 1.5 uL to .33uL A647
  • use old hybe mix (not hybe dilution mix)
  • stain in box without water (see how this goes)

Fly Work

  • flipped fly stocks
  • balancer Esc almost dead
  • Esc/CyO dead, but Esc/Gla still going strong (if stable?)
This entry was posted in Summaries and tagged , , . Bookmark the permalink.